# Elliptic Quantum Toroidal Algebra $U_{t_{1},t_{2},p}({\mathfrak{gl}}_{1,tor}),$ Vertex Operators and $L$-operators Hitoshi Konno and Andrey Smirnov

### v1

@@ 6.1 Type I vertex operator
$R_{T^{1/2}}(u_{1},u_{2};\mathfrak{z}\hbar^{c^{\perp(1)}+c^{\perp(2)}})=R_{T^{1/2}}(u_{1},u_{2};\mathfrak{z}).$ (6.8)
See Remark below (6.14). Note also that $\widehat{\mathrm{Stab}}_{{\mathfrak{C}},T^{1/2}}({{\boldsymbol{\alpha}}};\mathfrak{z})$ is depend on $u_{1},u_{2}$ only through the chamber $\mathfrak{C}:\ u_{1}\ll u_{2}$ essentially. Hence for any $a\in{\mathbb{C}}^{\times}$
@@
$\Phi^{(\mu^{\prime\prime})}_{\omega^{\prime\prime}}(u_{2})\Phi^{(\mu^{\prime})}_{\omega^{\prime}}(u_{1})=\sum_{{\boldsymbol{\lambda}}=(\lambda^{\prime},\lambda^{\prime\prime})\in({\cal P}^{2})_{n}}{R}_{T^{1/2}}(u_{1},u_{2};\mathfrak{z})^{\boldsymbol{\lambda}}_{{{\boldsymbol{\omega}}}}\ \Phi^{(\mu^{\prime})}_{\lambda^{\prime}}(u_{1})\Phi^{(\mu^{\prime\prime})}_{\lambda^{\prime\prime}}(u_{2}),$
where ${{\boldsymbol{\omega}}}=(\omega^{\prime},\omega^{\prime\prime})\in({\cal P}^{2})_{n}$.
@@
$\sum_{{\boldsymbol{\lambda}}=(\lambda^{\prime},\lambda^{\prime\prime})}\prod_{a\in\lambda^{\prime}}\left(\int_{0}^{\varphi^{\mu^{\prime}}_{\varsigma(a)}}d_{p}x_{a}\right)\prod_{a\in\lambda^{\prime\prime}}\left(\int_{0}^{\varphi^{\mu^{\prime\prime}}_{\varsigma(a)}}d_{p}x_{a}\right):\prod_{a\in{\boldsymbol{\lambda}}}x^{-}(x_{a}):\Phi_{\emptyset}(u_{2})\Phi_{\emptyset}(u_{1})$
×$\qquad\qquad\times\prod_{a,b\in{\boldsymbol{\lambda}}\atop\rho_{a}<\rho_{b}}<x^{-}(x_{a})x^{-}(x_{b})>^{sym}\ \widehat{\mathrm{Stab}}_{\mathfrak{C},T^{1/2}}({\boldsymbol{\lambda}};\mathfrak{z})\bar{R}_{T^{1/2}}(u_{1},u_{2};\mathfrak{z})^{\boldsymbol{\lambda}}_{{{\boldsymbol{\omega}}}}.$
Here we used (6.8). Note that $:\prod_{a\in{\boldsymbol{\lambda}}}x^{-}(x_{a}):$ and $\prod_{a,b\in{\boldsymbol{\lambda}}\atop\rho_{a}<\rho_{b}}<x^{-}(x_{a})x^{-}(x_{b})>^{sym}$ are symmetric in $x_{a}$ $(a\in{\boldsymbol{\lambda}})$ so that they are independent from a choice ${\boldsymbol{\lambda}}\in({\cal P}^{2})_{n}$. Evaluating the Jackson integral and using Proposition 5.1 and (6.5), one obtains
$\sum_{{\bf d}\in{\mathbb{N}}^{n}}\ \mathfrak{z}^{-\sum_{a\in\bullet}d_{a}}:\prod_{a\in\bullet}x^{-}(x_{a}):\bigr{|}_{x_{a}=\varphi^{\boldsymbol{\mu}}_{\varsigma(a)}p^{d_{a}}}\Phi_{\emptyset}(u_{2})\Phi_{\emptyset}(u_{1})\prod_{a,b\in\bullet\atop\rho_{a}<\rho_{b}}<x^{-}(x_{a})x^{-}(x_{b})>^{sym}\bigr{|}_{x_{a}=\varphi^{\boldsymbol{\mu}}_{\varsigma(a)}p^{d_{a}}}$
×$\hskip 85.35826pt\times\sum_{{\boldsymbol{\lambda}}}\widehat{\mathrm{Stab}}_{\mathfrak{C},T^{1/2}}({\boldsymbol{\lambda}};\mathfrak{z})\bigr{|}_{{\boldsymbol{\mu}}}\bar{R}_{T^{1/2}}(u_{1},u_{2};\mathfrak{z})^{\boldsymbol{\lambda}}_{{{\boldsymbol{\omega}}}}$
@@
×$\qquad\qquad\times\prod_{a,b\in\bullet\atop\rho_{a}<\rho_{b}}<x^{-}(x_{a})x^{-}(x_{b})>^{sym}\bigr{|}_{x_{a}=\varphi^{\boldsymbol{\mu}}_{\varsigma(a)}p^{d_{a}}}\times\widehat{\mathrm{Stab}}_{\overline{\mathfrak{C}},T^{1/2}}(\bar{{\boldsymbol{\omega}}};\mathfrak{z})\bigr{|}_{\bar{{\boldsymbol{\mu}}}}$
@@
Here $\bullet$ denotes any partition in $({\cal P}^{2})_{n}$. We also set $\{x_{a}=\varphi^{\boldsymbol{\mu}}_{\varsigma(a)}p^{d_{a}}\}:=\{x^{\prime}_{a}=\varphi^{\mu^{\prime}}_{\varsigma(a)}p^{d_{a}}\ (a\in\lambda^{\prime})\}\cup\{x^{\prime\prime}_{b}=\varphi^{\mu^{\prime\prime}}_{\varsigma(b)}p^{d_{b}}\ (b\in\lambda^{\prime\prime})\}$. ∎
6.1.1 Type II
@@
$\Psi^{*(\mu^{\prime\prime})}_{\lambda^{\prime\prime}}(u_{2})\Psi^{*(\mu^{\prime})}_{\lambda^{\prime}}(u_{1})=\prod_{a\in\lambda^{\prime\prime}}\left(\int_{0}^{\varphi^{\mu^{\prime\prime}}_{\varsigma(a)}}d_{p}x^{\prime\prime}_{a}\right)\prod_{a\in\lambda^{\prime}}\left(\int_{0}^{\varphi^{\mu^{\prime}}_{\varsigma(a)}}d_{p}x^{\prime}_{a}\right)\widehat{\mathrm{Stab}}^{*}_{\mathfrak{C},T^{1/2}_{opp}}((\lambda^{\prime},\lambda^{\prime\prime});\mathfrak{z}^{*-1}){}$
×$\hskip 56.9055pt\times:\prod_{a\in(\lambda^{\prime},\lambda^{\prime\prime})}x^{+}(x_{a}):\Psi^{*}_{\emptyset}(u_{2})\Psi^{*}_{\emptyset}(u_{1})\prod_{a,b\in(\lambda^{\prime},\lambda^{\prime\prime})\atop\rho_{a}<\rho_{b}}<x^{+}(x_{a})x^{+}(x_{b})>^{sym},$ (6.10)
where the chamber $\mathfrak{C}$ is given by $u_{1}\ll u_{2}$ with the stability condition $t_{1}/t_{2}>0$. Hence $\rho_{a}<\rho_{b}$ for $a\in\lambda^{\prime},b\in\lambda^{\prime\prime}$ due to (3.2). Similarly
$\Psi^{*(\mu^{\prime})}_{\lambda^{\prime}}(u_{1})\Psi^{*(\mu^{\prime\prime})}_{\lambda^{\prime\prime}}(u_{2})=\prod_{a\in\lambda^{\prime}}\left(\int_{0}^{\varphi^{\mu^{\prime}}_{\varsigma(a)}}d_{p}x^{\prime}_{a}\right)\prod_{a\in\lambda^{\prime\prime}}\left(\int_{0}^{\varphi^{\mu^{\prime\prime}}_{\varsigma(a)}}d_{p}x^{\prime\prime}_{a}\right)\widehat{\mathrm{Stab}}^{*}_{\overline{\mathfrak{C}},T^{1/2}_{opp}}((\lambda^{\prime\prime},\lambda^{\prime});\mathfrak{z}^{*-1}){}$
×$\hskip 56.9055pt\times:\prod_{a\in(\lambda^{\prime},\lambda^{\prime\prime})}x^{+}(x_{a}):\Psi^{*}_{\emptyset}(u_{1})\Psi^{*}_{\emptyset}(u_{2})\prod_{a,b\in(\lambda^{\prime},\lambda^{\prime\prime})\atop\rho_{a}<\rho_{b}}<x^{+}(x_{a})x^{+}(x_{b})>^{sym},$ (6.11)
where the chamber $\overline{\mathfrak{C}}$ is given by $u_{1}\gg u_{2}$ with the stability condition $t_{1}/t_{2}>0$. Hence $\rho_{a}<\rho_{b}$ for $a\in\lambda^{\prime\prime},b\in\lambda^{\prime}$.
Let us define $R^{*}_{T^{1/2}_{opp}}(u_{1},u_{2};\mathfrak{z}^{*-1})\in{\rm End}_{\mathbb{F}}({\cal F}^{(0,1)}_{u_{1}}\widetilde{\otimes}{\cal F}^{(0,1)}_{u_{2}})$ as the following transition matrix of the elliptic stable envelopes $\widehat{\mathrm{Stab}}^{*}_{{\mathfrak{C}},T^{1/2}_{opp}}(\bullet;\mathfrak{z}^{*-1})$. For ${\boldsymbol{\alpha}},{\boldsymbol{\beta}},{\boldsymbol{\gamma}}\in({\cal P}^{2})_{n}$,
@@
$\widehat{\mathrm{Stab}}^{*}_{\overline{\mathfrak{C}},T^{1/2}_{opp}}(\bar{{\boldsymbol{\alpha}}};\mathfrak{z}^{*-1})\bigr{|}_{\bar{{\boldsymbol{\gamma}}}}=\sum_{{\boldsymbol{\beta}}\in({\cal P}^{2})_{n}}\widehat{\mathrm{Stab}}^{*}_{{\mathfrak{C}},T^{1/2}_{opp}}({{\boldsymbol{\beta}}};\mathfrak{z}^{*-1})\bigr{|}_{{{\boldsymbol{\gamma}}}}\ \bar{R}^{*}_{T^{1/2}_{opp}}(u_{1},u_{2};\mathfrak{z}^{*-1})_{\boldsymbol{\alpha}}^{\boldsymbol{\beta}},$ (6.13)
@@
$\Psi^{*(\mu^{\prime})}_{\omega^{\prime}}(u_{1})\Psi^{*(\mu^{\prime\prime})}_{\omega^{\prime\prime}}(u_{2})=\sum_{{\boldsymbol{\lambda}}=(\lambda^{\prime},\lambda^{\prime\prime})\in({\cal P}^{2})_{n}}\Psi^{*(\mu^{\prime\prime})}_{\lambda^{\prime\prime}}(u_{2})\Psi^{*(\mu^{\prime})}_{\lambda^{\prime}}(u_{1}){R}^{*}_{T^{1/2}}(u_{1},u_{2};\mathfrak{z}^{*})_{{\boldsymbol{\lambda}}}^{\boldsymbol{\omega}},$
where ${\boldsymbol{\omega}}=(\omega^{\prime},\omega^{\prime\prime})\in({\cal P}^{2})_{n}$.
@@
$\Phi_{\lambda}(u)\Psi^{*}_{\mu}(v)=\chi(u/v)\Psi^{*}_{\mu}(v)\Phi_{\lambda}(u)\qquad\forall\lambda,\mu\in{\cal P}_{n},{}$
@@
$\Gamma(z;t_{1},t_{2})=\frac{(t_{1}t_{2}/z;t_{1},t_{2})_{\infty}}{(z;t_{1},t_{2})_{\infty}},\qquad(z;t_{1},t_{2})_{\infty}=\prod_{m_{1},m_{2}=0}^{\infty}(1-zt_{1}^{m_{1}}t_{2}^{m_{2}}).$
@@
Note $\gamma=\hbar^{1/2}$ and $p^{*}=p\hbar^{-1}$ in the level $(1,N)$ representation and the formula222 In this occasion let us correct a misprint in [22]. On page 431, the formula in (6.54) should read $\mu^{*}(v)=\frac{\chi(\frac{1}{2}-v)}{\chi(\frac{1}{2}+v)}\mu(v)|_{r\mapsto r^{*}}$.
@@
6.2 Dynamical Yang-Baxter equation
From the associativity of the composition of the type I and type II dual vertex operators, one can derive the dynamical Yang-Baxter equation for the elliptic $R$-matrices in (6.4) and (6.12).
For the type I vertex operator, using Proposition 6.1 for $\Phi^{(\mu^{\prime\prime\prime})}_{\omega^{\prime\prime\prime}}(u_{3})\Phi^{(\mu^{\prime\prime})}_{\omega^{\prime\prime}}(u_{2})\Phi^{(\mu^{\prime})}_{\omega^{\prime}}(u_{1})$ and assuming that $\Phi(u)$ is invertible, one obtains
$R^{(12)}_{T^{1/2}}(u_{1},u_{2};\mathfrak{z}\hbar^{c^{\perp(3)}})R^{(13)}_{T^{1/2}}(u_{1},u_{3};\mathfrak{z})R^{(23)}_{T^{1/2}}(u_{2},u_{3};\mathfrak{z}\hbar^{c^{\perp(1)}}){}$
$\qquad=R^{(23)}_{T^{1/2}}(u_{2},u_{3};\mathfrak{z})R^{(13)}_{T^{1/2}}(u_{1},u_{3};\mathfrak{z}\hbar^{c^{\perp(2)}})R^{(12)}_{T^{1/2}}(u_{1},u_{2};\mathfrak{z}),$ (6.17)

### v2

@@ 6.1. The type I vertex operator
$R_{T^{1/2}}(u_{1},u_{2};\mathfrak{z}\hbar^{c^{(1)}+c^{(2)}})=R_{T^{1/2}}(u_{1},u_{2};\mathfrak{z}).$ (6.8)
See Remark below (6.14). Note also that $\widehat{\mathrm{Stab}}_{{\mathfrak{C}},T^{1/2}}(\bullet;\mathfrak{z})$ depends on $u_{1},u_{2}$ only through the chamber $\mathfrak{C}:\ |u_{1}|>|u_{2}|$ essentially. Hence for any $a\in{\mathbb{C}}^{\times}$
@@
$\hskip-28.45274pt\Phi^{(\mu^{\prime\prime})}_{\omega^{\prime\prime}}(u_{2})\Phi^{(\mu^{\prime})}_{\omega^{\prime}}(u_{1})=\sum_{{\boldsymbol{\lambda}}=(\lambda^{\prime},\lambda^{\prime\prime})\in({\mathcal{P}}^{2})_{n}}{R}_{T^{1/2}}(u_{1},u_{2};\mathfrak{z})^{\boldsymbol{\lambda}}_{{{\boldsymbol{\omega}}}}\ \Phi^{(\mu^{\prime})}_{\lambda^{\prime}}(u_{1})\Phi^{(\mu^{\prime\prime})}_{\lambda^{\prime\prime}}(u_{2}),$
where ${{\boldsymbol{\omega}}}=(\omega^{\prime},\omega^{\prime\prime})\in({\mathcal{P}}^{2})_{n}$.
@@
$\hskip-28.45274pt\sum_{{\boldsymbol{\lambda}}=(\lambda^{\prime},\lambda^{\prime\prime})}\prod_{a\in\lambda^{\prime}}\left(\int_{0}^{\varphi^{\mu^{\prime}}_{\varsigma(a)}}d_{p}x_{a}\right)\prod_{a\in\lambda^{\prime\prime}}\left(\int_{0}^{\varphi^{\mu^{\prime\prime}}_{\varsigma(a)}}d_{p}x_{a}\right):\prod_{a\in{\boldsymbol{\lambda}}}x^{-}(x_{a}):{}$
$\qquad\times\Phi_{\emptyset}(u_{2})\Phi_{\emptyset}(u_{1})\prod_{a,b\in{\boldsymbol{\lambda}}\atop\rho_{a}<\rho_{b}}<x^{-}(x_{a})x^{-}(x_{b})>^{sym}{}$
$\qquad\times\ \widehat{\mathrm{Stab}}_{\mathfrak{C},T^{1/2}}({\boldsymbol{\lambda}};\mathfrak{z})\bar{R}_{T^{1/2}}(u_{1},u_{2};\mathfrak{z})^{\boldsymbol{\lambda}}_{{{\boldsymbol{\omega}}}}.$
Here we used (6.8). Note that $\prod_{a\in{\boldsymbol{\lambda}}}x^{-}(x_{a}):$ and $\prod_{a,b\in{\boldsymbol{\lambda}}\atop\rho_{a}<\rho_{b}}<x^{-}(x_{a})x^{-}(x_{b})>^{sym}$ are symmetric in $x_{a}$ $(a\in{\boldsymbol{\lambda}})$ so that they are independent from a choice ${\boldsymbol{\lambda}}\in({\mathcal{P}}^{2})_{n}$. Evaluating the Jackson integral and using Proposition 5.1 and (6.5), one obtains
$\hskip-28.45274pt\sum_{{\bf d}\in{\mathbb{N}}^{n}}\ \mathfrak{z}^{-\sum_{a\in\bullet}d_{a}}:\prod_{a\in\bullet}x^{-}(x_{a}):\bigr{|}_{x_{a}=\varphi^{\boldsymbol{\mu}}_{\varsigma(a)}p^{d_{a}}}\Phi_{\emptyset}(u_{2})\Phi_{\emptyset}(u_{1})$
$\quad\times\prod_{a,b\in\bullet\atop\rho_{a}<\rho_{b}}<x^{-}(x_{a})x^{-}(x_{b})>^{sym}\bigr{|}_{x_{a}=\varphi^{\boldsymbol{\mu}}_{\varsigma(a)}p^{d_{a}}}$
$\quad\times\sum_{{\boldsymbol{\lambda}}}\widehat{\mathrm{Stab}}_{\mathfrak{C},T^{1/2}}({\boldsymbol{\lambda}};\mathfrak{z})\bigr{|}_{{\boldsymbol{\mu}}}\bar{R}_{T^{1/2}}(u_{1},u_{2};\mathfrak{z})^{\boldsymbol{\lambda}}_{{{\boldsymbol{\omega}}}}$
@@
$\ \times\prod_{a,b\in\bullet\atop\rho_{a}<\rho_{b}}<x^{-}(x_{a})x^{-}(x_{b})>^{sym}\bigr{|}_{x_{a}=\varphi^{\boldsymbol{\mu}}_{\varsigma(a)}p^{d_{a}}}\times\widehat{\mathrm{Stab}}_{\overline{\mathfrak{C}},T^{1/2}}(\bar{{\boldsymbol{\omega}}};\mathfrak{z})\bigr{|}_{\bar{{\boldsymbol{\mu}}}}$
@@
Here $\bullet$ denotes any partition in $({\mathcal{P}}^{2})_{n}$. We also set $\{x_{a}=\varphi^{\boldsymbol{\mu}}_{\varsigma(a)}p^{d_{a}}\}:=\{x^{\prime}_{a}=\varphi^{\mu^{\prime}}_{\varsigma(a)}p^{d_{a}}\ (a\in\lambda^{\prime})\}\cup\{x^{\prime\prime}_{b}=\varphi^{\mu^{\prime\prime}}_{\varsigma(b)}p^{d_{b}}\ (b\in\lambda^{\prime\prime})\}$. ∎
6.2. The type II dual vertex operator
@@
$\hskip 11.38109pt\Psi^{*(\mu^{\prime\prime})}_{\lambda^{\prime\prime}}(u_{2})\Psi^{*(\mu^{\prime})}_{\lambda^{\prime}}(u_{1})$ (6.10)
$\hskip-19.91692pt=\prod_{a\in\lambda^{\prime\prime}}\left(\int_{0}^{\varphi^{\mu^{\prime\prime}}_{\varsigma(a)}}d_{p}x^{\prime\prime}_{a}\right)\prod_{a\in\lambda^{\prime}}\left(\int_{0}^{\varphi^{\mu^{\prime}}_{\varsigma(a)}}d_{p}x^{\prime}_{a}\right)\widehat{\mathrm{Stab}}^{*}_{\overline{\mathfrak{C}},T^{1/2}_{opp}}((\lambda^{\prime},\lambda^{\prime\prime});\mathfrak{z}^{*-1}){}$
$\hskip-19.91692pt\quad\times:\hskip-8.53581pt\prod_{a\in(\lambda^{\prime},\lambda^{\prime\prime})}x^{+}(x_{a}):\Psi^{*}_{\emptyset}(u_{2})\Psi^{*}_{\emptyset}(u_{1})\prod_{a,b\in(\lambda^{\prime},\lambda^{\prime\prime})\atop\rho_{a}>\rho_{b}}\hskip-11.38109pt<x^{+}(x_{a})x^{+}(x_{b})>^{sym},{}$
where the chamber $\overline{\mathfrak{C}}$ is given by $|u_{1}|<|u_{2}|$ with the stability condition $t_{1}/t_{2}>0$. Similarly
$\hskip 11.38109pt\Psi^{*(\mu^{\prime})}_{\lambda^{\prime}}(u_{1})\Psi^{*(\mu^{\prime\prime})}_{\lambda^{\prime\prime}}(u_{2})$ (6.11)
$\hskip-19.91692pt=\prod_{a\in\lambda^{\prime}}\left(\int_{0}^{\varphi^{\mu^{\prime}}_{\varsigma(a)}}d_{p}x^{\prime}_{a}\right)\prod_{a\in\lambda^{\prime\prime}}\left(\int_{0}^{\varphi^{\mu^{\prime\prime}}_{\varsigma(a)}}d_{p}x^{\prime\prime}_{a}\right)\widehat{\mathrm{Stab}}^{*}_{{\mathfrak{C}},T^{1/2}_{opp}}((\lambda^{\prime\prime},\lambda^{\prime});\mathfrak{z}^{*-1}){}$
$\hskip-19.91684pt\quad\times:\hskip-8.53581pt\prod_{a\in(\lambda^{\prime},\lambda^{\prime\prime})}x^{+}(x_{a}):\Psi^{*}_{\emptyset}(u_{1})\Psi^{*}_{\emptyset}(u_{2})\prod_{a,b\in(\lambda^{\prime},\lambda^{\prime\prime})\atop\rho_{a}>\rho_{b}}\hskip-11.38109pt<x^{+}(x_{a})x^{+}(x_{b})>^{sym},{}$
where the chamber $\mathfrak{C}$} is given by $|u_{1}|>|u_{2}|$ with the stability condition $t_{1}/t_{2}>0$.
Let us define $R^{*}_{T^{1/2}_{opp}}(u_{1},u_{2};\mathfrak{z}^{*-1})\in{\rm End}_{\mathbb{F}}({\mathcal{F}}^{(0,-1)}_{u_{1}}\widetilde{\otimes}{\mathcal{F}}^{(0,-1)}_{u_{2}})$ as the following transition matrix of the elliptic stable envelopes $\widehat{\mathrm{Stab}}^{*}_{{\mathfrak{C}},T^{1/2}_{opp}}(\bullet;\mathfrak{z}^{*-1})$. For ${\boldsymbol{\alpha}},{\boldsymbol{\beta}},{\boldsymbol{\gamma}}\in({\mathcal{P}}^{2})_{n}$,
@@
$\widehat{\mathrm{Stab}}^{*}_{{\mathfrak{C}},T^{1/2}_{opp}}(\bar{{\boldsymbol{\alpha}}};\mathfrak{z}^{*-1})\bigr{|}_{\bar{{\boldsymbol{\gamma}}}}$ (6.13)
$=\sum_{{\boldsymbol{\beta}}\in({\mathcal{P}}^{2})_{n}}\widehat{\mathrm{Stab}}^{*}_{\overline{\mathfrak{C}},T^{1/2}_{opp}}({{\boldsymbol{\beta}}};\mathfrak{z}^{*-1})\bigr{|}_{{{\boldsymbol{\gamma}}}}\ \bar{R}^{*}_{T^{1/2}_{opp}}(u_{1},u_{2};\mathfrak{z}^{*-1})_{\boldsymbol{\alpha}}^{\boldsymbol{\beta}},{}$
@@
$\hskip-28.45274pt\Psi^{*(\mu^{\prime})}_{\omega^{\prime}}(u_{1})\Psi^{*(\mu^{\prime\prime})}_{\omega^{\prime\prime}}(u_{2})=\hskip-8.5359pt\sum_{{\boldsymbol{\lambda}}=(\lambda^{\prime},\lambda^{\prime\prime})\in({\mathcal{P}}^{2})_{n}}\hskip-8.5359pt\Psi^{*(\mu^{\prime\prime})}_{\lambda^{\prime\prime}}(u_{2})\Psi^{*(\mu^{\prime})}_{\lambda^{\prime}}(u_{1}){R}^{*}_{T^{1/2}}(u_{1},u_{2};\mathfrak{z}^{*})_{{\boldsymbol{\lambda}}}^{\boldsymbol{\omega}},$
where ${\boldsymbol{\omega}}=(\omega^{\prime},\omega^{\prime\prime})\in({\mathcal{P}}^{2})_{n}$.
@@
$\Phi_{\lambda}(u)\Psi^{*}_{\mu}(v)=\chi(u/v)\Psi^{*}_{\mu}(v)\Phi_{\lambda}(u)\qquad\forall\lambda,\mu\in{\mathcal{P}}_{n},{}$
@@
$\hskip-28.45274pt\Gamma(z;t_{1},t_{2})=\frac{(t_{1}t_{2}/z;t_{1},t_{2})_{\infty}}{(z;t_{1},t_{2})_{\infty}},\quad(z;t_{1},t_{2})_{\infty}=\prod_{m_{1},m_{2}=0}^{\infty}(1-zt_{1}^{m_{1}}t_{2}^{m_{2}}).$
@@
Note $\gamma=\hbar^{1/2}$ and $p^{*}=p\hbar^{-1}$ in the level $(1,N)$ representation and the formula333 In this occasion let us correct a misprint in [22]. On page 431, the formula in (6.54) should read $\mu^{*}(v)=\frac{\chi(\frac{1}{2}-v)}{\chi(\frac{1}{2}+v)}\mu(v)|_{r\mapsto r^{*}}$.
@@
6.3. Dynamical Yang-Baxter equation
From the associativity of the composition of the type I and the type II dual vertex operators, one can derive the dynamical Yang-Baxter equation for the elliptic $R$-matrices in (6.4) and (6.12).
For the type I vertex operator, using Proposition 6.1 for
$\Phi^{(\mu^{\prime\prime\prime})}_{\omega^{\prime\prime\prime}}(u_{3})\Phi^{(\mu^{\prime\prime})}_{\omega^{\prime\prime}}(u_{2})\Phi^{(\mu^{\prime})}_{\omega^{\prime}}(u_{1})$ and assuming that $\Phi(u)$ is invertible, one obtains
$R^{(12)}_{T^{1/2}}(u_{1},u_{2};\mathfrak{z}\hbar^{-c^{(3)}})R^{(13)}_{T^{1/2}}(u_{1},u_{3};\mathfrak{z})R^{(23)}_{T^{1/2}}(u_{2},u_{3};\mathfrak{z}\hbar^{-c^{(1)}})$ (6.17)
$\quad=R^{(23)}_{T^{1/2}}(u_{2},u_{3};\mathfrak{z})R^{(13)}_{T^{1/2}}(u_{1},u_{3};\mathfrak{z}\hbar^{-c^{(2)}})R^{(12)}_{T^{1/2}}(u_{1},u_{2};\mathfrak{z}),{}$